OpenStudy (anonymous):

The area of a rectangular field is equal to 300 square meters. Its perimeter is equal to 70 meters. Find the length and width of this rectangle.

Mehek (mehek14):

L: Length of the Rectangle W: Width of Rectangle Area of Rectangular : Length x Width = L*W =300---(i) Perimeter = 2(L+W)=70 or 2L + 2W = 70 (L+W)=35 L=35-W Put this value in Eq (i), you will get following equation : x2-35x+300 = 0 After solving this equation, L=20 and W=15

OpenStudy (anonymous):

okay thx @Mehek14

Mehek (mehek14):

yw @shifuyanli